Titration lab

State the purpose of this experiment.​

Describe the function of the phenolphthalein in this experiment. How would this experiment be different if you had forgotten to add the phenolphthalein

State evidence from the lab that the endpoint was reached.

It takes 75ml of a 2.5M HCl solution to neutralize 55ml of a base of unknown concentration. Calculate the concentration of an unknown basic solution.

If the actual concentration of the HCl was 0.33M what is your percent error?

Environmental studies usually involve an analysis of precipitation and its response to pollution. To quantify the degree of contamination in natural rain water or snow, titration is used. The process is quick and results are reliable. Since most titration processes do not require expensive or specialized equipment, the test can be performed often and in different areas with relatively little effort.
a. A 1000.0 mL sample of lake water is titrated using 0.100 mL of a 0.100 M base solution. What is the molarity of the acid in the lake water?
b. Based on the molarity of the acid calculated above, what is the pH of the lake water?
Answer #1

Purpose is to get the concentration of an unknown using a known solution of known concentration which can react with each other.Titration can help determine unknown concentration and even the identity of unknown.

We know that V1N1=V2N2.

75×2.5=55×y

y=3.41 M

If it were 0.33 M, it would have been=75×0.33/55=0.45 M

%error=3.41-0.45/0.45×100=657%

A. 1000×y=0.1×0.1

y=0.00001 M acid

b. pH=-logH+=5